# Parameter-Efficient Fine-Tuning of LLaMA for the Clinical Domain

### v1

@@ 5 Conclusions
In this study, we propose a two-step PEFT framework. We introduce Clinical LLaMA-LoRA, a LoRA Hu et al. (2022) adapter built upon LLaMA Touvron et al. (2023). Then, we introduce Downstream LLaMA-LoRA, a task-specific adapter that is trained on top of the pretrained Clinical LLaMA-LoRA. The fusion of the two adapters achieves state-of-the-art performance with an AUROC score of 72.81% macro-averaged across all clinical NLP downstream tasks, which represents a 3.22% improvement over the previous best-performing model. Our proposed framework achieves improvement in performance while reducing the computational requirements, which is suited for clinical settings that are often constrained by their computational power.
We also find that the LoRA-equipped BlueBERT model achieves a considerable improvement of macro-averaged AUROC score over the full fine-tuning (71.56% compared to 69.59%), with notable improvements in mortality and length-of-stay prediction. These findings further highlight the potential to achieve strong performance without extensive computational resources.
Future works may explore developing a schema to address various real-world use cases, building upon the findings of this study. Such a schema would use multiple Downstream LLaMA-LoRA adapters tailored for different use cases while leveraging the pretrained LLM and Clinical LLaMA-LoRA as the foundation. This solution would also be suited for use cases which rely on private data commonly encountered in care provider settings.
@@
This study presents a two-step PEFT framework aimed at effectively adapting LLMs to diverse clinical downstream applications. However, the evaluation of our model was restricted to MIMIC-based datasets, which are constrained to English and obtained exclusively within the Commonwealth of Massachusetts, United States of America. Consequently, despite the promising efficacy demonstrated by our proposed method, it would have been advantageous to directly assess its performance across diverse hospital systems spanning various geographical locations and languages. This would enable a more comprehensive understanding of its applicability and generalizability. However, it is essential to acknowledge that conducting such an analysis would require working within a trusted research environment and obtaining the necessary permissions to access the relevant datasets.
@@
Despite the demonstrated performance improvements, the proposed model may still be susceptible to spurious correlations. Predicting patient outcomes solely based on clinical notes presents significant challenges due to the other factors that may not be captured within those notes. For instance, the length of patient’s in-hospital stay is not solely correlated with their diagnoses and disease progression. Factors such as the patient’s insurance status, which is not typically mentioned in clinical notes, can severely impact the duration of a patient’s stay. Therefore, we encourage end users of such clinical LLMs to consider additional measures to ensure predictions that reflect a holistic view of the patient’s situation, instead of relying solely on the predictions of LLMs.
@@ Acknowledgements
AG was supported by the United Kingdom Research and Innovation (grant EP/S02431X/1), UKRI Centre for Doctoral Training in Biomedical AI at the University of Edinburgh, School of Informatics. PM was partially funded by the European Union’s Horizon 2020 research and innovation programme under grant agreement no. 875160, ELIAI (The Edinburgh Laboratory for Integrated Artificial Intelligence) EPSRC (grant no. EP/W002876/1), an industry grant from Cisco, and a donation from Accenture LLP; and is grateful to NVIDIA for the GPU donations. BA was partially funded by by Legal and General PLC as part of the Advanced Care Research Centre and by the Artificial Intelligence and Multimorbidity: Clustering in Individuals, Space and Clinical Context (AIM-CISC) grant NIHR202639. For the purpose of open access, AG has applied a creative commons attribution (CC BY) licence to any author-accepted manuscript version arising.
@@ A.1 Fixed Model Hyperparameters
Maximum sequence length 128
@@ B.1 Fixed Model Hyperparameters
Maximum sequence length 128
@@ B.2 PEFT Hyperparameters Optimisation Search Space
Appendix C Training Configurations

### v3

@@ 5 Conclusions
In this study, we propose a two-step PEFT framework. We introduce Clinical LLaMA-LoRA, a LoRA Hu et al. (2022) adapter built upon LLaMA Touvron et al. (2023). Then, we introduce Downstream LLaMA-LoRA, a task-specific adapter that is trained on top of the pretrained Clinical LLaMA-LoRA. The fusion of the two adapters achieves an AUROC score of 76.07% macro-averaged across all clinical NLP downstream tasks, which represents a 3.37% improvement over the best-performing clinical LLM. Our proposed framework achieves improvement in performance while reducing the computational requirements, which is suited for clinical settings that are often constrained by their computational power.
@@
This study presents a two-step PEFT framework aimed at effectively adapting LLMs to diverse clinical downstream applications. However, the evaluation of our model was restricted to MIMIC-based datasets, which are constrained to English and obtained exclusively within the Commonwealth of Massachusetts, United States of America. Consequently, despite the promising efficacy demonstrated by our proposed method, it would have been advantageous to directly assess its performance across diverse hospital systems spanning other geographical locations and languages. This would enable a more comprehensive understanding of its applicability and generalizability. However, it is essential to acknowledge that conducting such an analysis would require working within a trusted research environment and obtaining the necessary permissions to access the relevant datasets.
@@
Despite the demonstrated performance improvements, the proposed model may still be susceptible to spurious correlations. Predicting patient outcomes solely based on clinical notes presents significant challenges due to the other factors that may not be captured within those notes. For instance, the length of a patient’s in-hospital stay is not solely correlated with their diagnoses and disease progression. Factors such as the patient’s insurance status, which is not typically mentioned in clinical notes, can severely impact the duration of a patient’s stay. Therefore, we encourage end users of such clinical LLMs to consider additional measures to ensure predictions that reflect a holistic view of the patient’s situation, instead of relying solely on the predictions of LLMs.
@@ Acknowledgements
APG was supported by the United Kingdom Research and Innovation (grant EP/S02431X/1), UKRI Centre for Doctoral Training in Biomedical AI at the University of Edinburgh, School of Informatics. PM was partially funded by the European Union’s Horizon 2020 research and innovation programme under grant agreement no. 875160, ELIAI (The Edinburgh Laboratory for Integrated Artificial Intelligence) EPSRC (grant no. EP/W002876/1), an industry grant from Cisco, and a donation from Accenture LLP; and is grateful to NVIDIA for the GPU donations. BA was partially funded by Legal and General PLC as part of the Advanced Care Research Centre and by the Artificial Intelligence and Multimorbidity: Clustering in Individuals, Space and Clinical Context (AIM-CISC) grant NIHR202639. For the purpose of open access, the authors have applied a Creative Commons attribution (CC BY) licence to any author-accepted manuscript version arising. This work was supported by the Edinburgh International Data Facility (EIDF) and the Data-Driven Innovation Programme at the University of Edinburgh.
We also thank Yahan Li, Ajitha Rajan, and Bruce Guthrie for their feedback on the manuscript.
@@ A.1 Fixed Model Hyperparameters
Maximum sequence length 512
@@ B.1 Fixed Model Hyperparameters
Maximum sequence length 512
@@ B.2 PEFT Hyperparameters Optimisation Search Space
Appendix C Comparison with BEEP Naik et al. (2022)
Model PMV MOR LOS Avg
BEEP 59.43 84.65 72.71 72.26
Our method 53.81 83.02 73.26 70.03
Table 8: AUROC scores in a subset of the clinical downstream document classification tasks. The macro-averaged AUROC score is calculated by taking the average of AUROC scores across this subset of tasks. The row in italic indicates the model variant with the highest macro-averaged AUROC.
We compared our method with the state-of-the-art clinical outcome prediction model, BEEP Naik et al. (2022), which leverages a retrieval augmentation technique to enhance the predictive capabilities of clinical language models. A small caveat is that BEEP focused on three downstream tasks: prolonged mechanical ventilation, mortality, and length of stay predictions. We selected the best-performing solution from BEEP, UmlsBERT with weighted voting retrieval augmentation, based on the averaged AUROC score to compare with our solution. While BEEP outperforms our approach, particularly in the prediction of PMV, it is crucial to emphasise that our method achieves its predictions without relying on retrieval augmentation. Future work may explore using retrieval augmentation on top of our proposed method.
Appendix D Training Configurations
@@
Appendix E Artefacts
The pretrained baseline models including BioClinicalBERT Alsentzer et al. (2019), BlueBERT Peng et al. (2019), and CORe van Aken et al. (2021) were released under the Creative Commons designation CC0 1.0 Universal license, whereas UmlsBERT Michalopoulos et al. (2021) was released under the MIT license. LLaMA Touvron et al. (2023) was released under a noncommercial license.
MIMIC-III and MIMIC-IV dataset was released under the PhysioNet Credentialed Health Data License 1.5.0 and can only be accessed after one finishes the CITI Data or Specimens Only Research training [^3] https://physionet.org/about/citi-course/.